OpenStudy (anonymous):

A square is inscribed inside a circle. The square's vertices have coordinates R(0, 4), S(6, 2), T(4, -4), and U(-2, -2). Find the equation of the circle. (x-?)^2+y^2=?

OpenStudy (anonymous):

guess we need the center and radius. let me draw a picture. center will be halfway between the diagonal of the square, and radius will be half the length of the diagonal. i am guessing that you can find both of those, but if not let me know

OpenStudy (vishweshshrimali5):

satellite can u help me too please

OpenStudy (anonymous):



OpenStudy (anonymous):

we can use \((0,4)\) and \((4,-4)\) as the two points we need the midpoint and the distance. let me know if you need help finding those

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

to find the midpoint, take the average of the coordinates \[(\frac{0+4}{2},\frac{4-4}{2})\] \[(2,0)\] is the center

OpenStudy (anonymous):

now for the radius squared, find the square of the distance between \((2,0)\) and \((0,4)\) by pythagoras. it is \[2^2+4^2=4+16=20\] now you have both the center \((2,0)\) and \(r^2=20\) so you can write the equation